1) two hetrozygoys blond/brown hair parents have 5 children. what is the probabilty that:

a) the first 3 children will be blond

b) three children will have blond hair

c)at least three children will have blond hair   

explain please ...
Answer #1

1. The parents are heterozygous, let their genotypes be Bb and Bb.

Children can have the four possible genotypes- BB,Bb,Bb,bb

Let the genotype of blonde hair be bb.Then probability of a child having bb genotype that is blonde hair is 1/4

a).Therefore probability that the first 3 children will be blonde is 1/4x1/4x1/4=1/64

b). Total number of chidren =5

If 3 chidren have blonde hair then 2 chidren will have brown hair.

Therefore probability that three children will have blond hair= 3/5

c). probability that atleast three children will have blond hair=1-2/5